A sequence begins 2,14,26,38,50 and each term is 12 more than the last. Whats the equation?
Answer by Alan3354(69443)

A sequence begins 2,14,26,38,50 and each term is 12 more than the last. Whats the equation?
-----------
term(n+1) = term(n) + 12
-----
term(n) = 2 + 12n, n = 0,1,2,3...